The proclist package^{∗}

Andrew D. Lewis

1998/04/03

1 Using the proclist macros

The package proclist (proclamation lists) primarily provides an environment

proclist proclist which is suitable for enumeration within theorem environments. I find the enumerate environment leaves to much space between items and I do not like the use of arabic numbering, etc. The environment works exactly like the enumerate environment, however. It just looks different.

I have also added some referencing macros to do some common tasks in a

\procref consistent manner. First is which you can use to refer to any item you labelled within a proclist environment. It will surround the expression in parentheses. I often want to refer to the proclamation number at the same time as I refer to the

\procdblref list item. To do this use the syntax

\pldblref{label for theorem}{label for list item}

\enumdblref This will give you a nice looking reference. I have also added a macro \enumdblref which does the same thing for the usual enumerate environment.

2 Implementation

The environment proclist is a thinly disguised version of enumerate. Not too much to describe here. One has to find a good way to make the references look not too shabby. I hope you like what I did.

1 ⟨∗package⟩

2

 3 \newcount\@ADL@plistdepth\@ADL@plistdepth\z@
 4 \@definecounter{plisti}
 5 \@definecounter{plistii}
 6 \@definecounter{plistiii}
 7 \@definecounter{plistiv}
 8 \newenvironment{proclist}{\ifnum \@ADL@plistdepth >3 \@toodeep\else
 9 \advance\@ADL@plistdepth \@ne

∗This file has version number v1.0b, last revised 1998/04/03.

10 \edef\@ADL@plistctr{plist\romannumeral\the\@ADL@plistdepth}\list
11 {\csname label\@ADL@plistctr\endcsname}{\usecounter{\@ADL@plistctr}
12 \topsep 3\p@ plus \p@ minus \p@\parsep 1.5\p@ plus \p@ minus .5\p@
13 \itemsep\parsep\partopsep \p@ plus \p@ minus \p@
14 \def\makelabel##1{\hss\llap{##1}}}\fi}
15 {\endlist}

16

17 \renewcommand{\theplisti}{\roman{plisti}}
18 \renewcommand{\theplistii}{\alph{plistii}}
19 \renewcommand{\theplistiii}{\Roman{plistiii}}
20 \renewcommand{\theplistiv}{\Alph{plistiv}}
21 \newcommand{\labelplisti}{(\theplisti)}
22 \newcommand{\labelplistii}{(\theplistii)}
23 \newcommand{\labelplistiii}{\theplistiii.}
24 \newcommand{\labelplistiv}{\theplistiv.}
25 \renewcommand\p@plistii{\theplisti\,}
26 \renewcommand\p@plistiii{\theplisti\,\theplistii\,}
27 \renewcommand\p@plistiv{\p@plistiii\theplistiii\,}

Here are the referencing commands.

28

29 \newcommand\plref[1]{\mbox{(\ref{#1})}}
30 \newcommand\pldblref[2]{\mbox{{\@ADL@plreffont\ref{#1}(\ref{#2})}}}
31 \newcommand\enumdblref[2]{\mbox{{\@ADL@enumreffont\ref{#1}%
32 \textendash\ref{#2}}}}
33 \def\@ADL@plreffont{\relax}
34 \def\@ADL@enumreffont{\relax}

35

36 ⟨/package⟩